Problem # 1: Solve -3x - 4y + 2z = -6
Problem # 2: Solve 4x - 6y + 3z = 8
Problem # 3: Solve y^2 - 16/2y + 6 divided by y -4/ y + 3
Problem # 4: Which of the following ordered pairs is a solution of y =x - 8:
a(20,13) b(12,20) c(12,21) d(20,12)

Answer by nycsub_teacher(90)

I will only do one of these questions.
Problem # 4: Which of the following ordered pairs is a solution of y =x - 8:
a(20,13) b(12,20) c(12,21) d(20,12)
We need to find a point that will produce the same answer on BOTH sides after we plug and chug.
How about point d?
You were given the linear equation: y = x - 8
Let x = 20
Let y = 12
12 = 20 - 8
12 = 12...IT CHECKS!!
The answer to question 4 is choice d.